# Can You Trust LLM Judgments? Reliability of LLM-as-a-Judge

### v1

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have experienced rapid advancements and widespread adoption. Such growth has been embraced widely, with applications ranging from marketing to biotechnology to music (Gozalo-Brizuela and Garrido-Merchán, 2023). As reliance upon LLMs grows, ensuring the quality and reliability of their outputs becomes crucial.
@@
A singular output from an LLM is the result of a draw from a probabilistic distribution (Vaswani, 2017). Parameters that vary the resulting draw, namely temperature and top-k, are increased to obtain more novel or "creative" results - a desirable quality of LLMs (Peeperkorn et al., 2024). However, this inherent stochastic nature raises concerns about the reliability of individual outputs. Given that an output from an LLM is only a single instantiation of the model, reliability of LLM output is of utmost importance (Mittelstadt et al., 2023). Consider a scenario where an LLM is tasked with evaluating the quality of another LLM that’s handling social media moderation. In this high-stakes context, ensuring the reliability of LLM judgments is crucial to user safety. If an LLM’s judgment is influenced by random chance rather than its underlying capabilities, it could lead to serious consequences, such as the proliferation of harmful content.
@@
Many research works aim to circumvent this issue by setting a fixed instantiation and using deterministic settings for the temperature and top-k parameters (Ouyang et al., 2023; Wei et al., 2024; Atil et al., 2024). These studies argue that if an LLM consistently produces the same output under these conditions, it can be considered reliable. However, consistent replication does not guarantee the reliability of the generated text.
Even with deterministic settings, a single LLM output remains a sample from the model’s probability distribution, subject to inherent randomness. This results in "fixed randomness," which can lead to significant limitations (Hellrich and Hahn, 2016). Consider our prior example of an LLM tasked with evaluating the quality of another LLM handling social media moderation. If the judging LLM fails to detect poor performance in a single instantiation due to randomness, grave repercussions can result. By assessing the reliability of LLM judgments, we can determine when additional human oversight is necessary to ensure the safety and accuracy of LLM-powered systems, especially as these systems evolve over time.
@@
While some research has focused on determining LLM reliability by comparing outputs to ground truth labels (Zhou et al., 2024; Fu et al., 2023a), this approach is often limited by the availability of accurate and comprehensive ground truth data. As LLMs are more readily used in more complex and subjective contexts, such as in the scenario of an LLM tasked with evaluating the quality of another LLM handling social media moderation, existence of a ground truth response becomes increasingly impractical. Other studies have explored inter-rater reliability (Kollitsch et al., 2024; Wang et al., 2024), but this metric is not suitable for evaluating LLM judges, as it assumes multiple human evaluators rather than a single, potentially biased, LLM model. Additionally, inter-rater reliability does not take into account the additional nuance of the LLM judge being a model itself.
The unique characteristics of LLMs as evaluators necessitate a more nuanced approach to assessing their reliability. We need to consider factors such as the consistency of their judgments, their sensitivity to different prompts and datasets, and their ability to generalize to new domains. A more robust framework for assessing LLM-based evaluation is essential to ensure quality and fairness in AI-driven decision-making. A key component of this framework is a measure of the internal consistency reliability of LLM judgments.
Internal consistency reliability, a well-established statistical concept with a rich literature, can be effectively measured using metrics like Cronbach’s alpha and McDonald’s omega (Cronbach, 1951; McDonald, 2013; Malkewitz et al., 2023). These metrics can effectively quantify reliability of a model’s judgments by assessing the consistency of its responses to different prompts or tasks.
We introduce an innovative framework for evaluating the reliability of LLMs as judges. We illustrate that McDonald’s omega can effectively quantify the internal consistency of LLM judgments across replications (different runs where only the instantiation changes). Applying this to both single-turn and multi-turn tasks across varying temperatures, we investigate the impact of LLM-as-a-judge reliability. Further, we demonstrate the implications of reliability on downstream applications.  [^2] The code for the framework and analyses can be accessed at https://github.com/kaylaschroeder/llm_reliability.
@@ 2 Related Work
Internal consistency reliability itself, especially Cronbach’s alpha, has been heavily utilized as an assessor of question quality in LLM contexts (Shang et al., 2024; Chan et al., 2023b; Biri et al., 2023). Bhandari et al. (2024) uses Cronbach’s alpha to compare LLM-generated questions with human-authored ones; Shi et al. (2023) uses Cronbach’s alpha to assess the reliability of questions designed to evaluate LLMs.
@@
The concept of reliability appears widely in the LLM literature, but is not consistently defined. Some studies equate reliability with human-model agreement, requiring costly human evaluations (Sun et al., 2024; Zhang et al., 2024). Others rely on accuracy metrics, which necessitate ground truth labels (Zheng et al., 2024; Macherla et al., 2023; Wei et al., 2024), a requirement that proves increasingly unrealistic as the judgment task increases in complexity and subjectivity. Semantic consistency has also been proposed as a reliability measure (Govil et al., 2024; Raj et al., 2023), despite its lack of theoretical foundation and potential pitfalls, as detailed in Schroeder and Wood-Doughty (2024). Many studies also use the term "reliability" without rigorous definition (Tan et al., 2024; Liao et al., 2023; Gupta et al., 2024).
@@
Currently, the LLM-as-a-judge literature predominantly relies upon accuracy and bias metrics to assess reliability (Chen et al., 2024; Wei et al., 2024; Ye et al., 2024). Some researchers advocate for combining responses from different LLM judgment models to improve reliability (Gu et al., 2024; Patel et al., 2024). To quantify reliability, human judgments are often used as a benchmark, with methods relying upon agreement with human evaluation and correlation with human-annotated error scores (Jung et al., 2024; Fu et al., 2023b; Dong et al., 2024). Our work pioneers the investigation of internal consistency reliability in the LLM-as-a-judge paradigm, addressing a critical gap in the literature.
3 Internal Consistency Reliability
@@
Internal consistency reliability is a cornerstone of psychometric analysis and has been extensively studied. Cronbach’s alpha, introduced by Cronbach (1951), has been widely used to assess internal consistency, but relies on several assumptions that may not always hold (Malkewitz et al., 2023; Agbo, 2010). As proposed by McDonald (2013), McDonald’s omega addresses these limitations by accounting for a broader range of factor structures. Ultimately, omega provides a more robust measure of internal consistency and has been shown to be more robust to deviations from assumptions than alpha (Stensen and Lydersen, 2022). In this work, we adopt McDonald’s omega as our primary metric, recognizing its superior quantification of reliability.
@@ 3.1 McDonald’s Omega
where (in this setting) $n$ is the number of replications, $\lambda_{i}$ is the factor loading of the $i$th replication, $\theta_{ii}$ is the standardized error variance of the $i$th replication, and $\theta_{ij}$ is the covariance between replications $i$ and $j$ (McDonald, 2013). A replication in this context refers to an LLM judgment output that differs only in its instantiation from other replications for the same judgment prompt. Omega relies upon the common multidimensional methodology of factor analysis, of which the factor loadings are derived (Kline, 2014). Factor loadings are statistical measures derived from factor analysis. They quantify the strength of the relationship between each replication’s result and the underlying latent trait, which is the unobserved true judgment that the replications are intended to measure. In simpler terms, factor loadings indicate how much each replication contributes to our understanding of the overall judgment being measured. A higher factor loading suggests that the replication is a strong indicator of the latent trait.
@@
    Additive measurement error: Each replication’s values are subject to random errors that are additive to the true score.
@@
    Uncorrelated errors: The errors associated with different replications are independent of each other.
@@
    Single latent trait: Each replication measures a single underlying trait.
Assumption 1, requiring additive measurement error, is a common assumption in statistical test theory, especially for items with similar response formats, making it a reasonable assumption in our context (Alagumalai and Curtis, 2005; Wadkar et al., 2016; Liu et al., 2010). Additionally, assumptions 2 and 3 are trivial in our setting as each replication is independent so errors are guaranteed to be uncorrelated, and each replication result is an outputted response to a singular prompt so it logically follows that each replication measures a single underlying construct.
@@ 4 Experiments and Data
We evaluate the reliability of the LLM-as-a-judge paradigm across diverse question formats and difficulty levels. Using LLM responses to benchmark questions, judgment LLMs are prompted repeatedly to select the "best" response based on factors like accuracy, utility, and relevance. We ran each judgment prompt 100 times, varying only the instantiation while holding other factors constant. Additionally, we varied the temperature parameter to further explore the influence of fixed randomness on LLM outputs. All judgment results were then assessed to uncover reliability by applying McDonald’s omega.
To obtain data for judgments, we utilize the widely accepted benchmarks BIG-Bench Hard (BBH) from Suzgun et al. (2022), SQuAD from Rajpurkar (2016), and MT-Bench from Zheng et al. (2023). These benchmarks offer two key advantages. First, they present diverse question formats: BBH and SQuAD questions are single-turn with specific target answers, while MT-Bench questions are multi-turn and open-ended. Second, they provide varying levels of difficulty, as evidenced by the significantly different accuracy rates of LLMs on BBH and SQuAD questions shown in Table 3. This diversity allows us to assess judgments under a range of conditions.
@@
We first randomly selected one question from each category within each benchmark, for a total of 55 questions. Next, we generated five responses for each question using Chain-of-Thought prompting, as illustrated in Figure 1. The variation in prompting by benchmark is a result of the structuring of each benchmark; BBH responses require a single question, SQuAD responses require a single question as well as context surrounding the question, and MT-Bench requires a two-turn dialogue. Responses to questions are generated from five highly popular open-source LLMs: LLaMA-3-8B, Vicuna-7B-v1.5, Gemma-7B, Phi-2, and Falcon-7b (Dubey et al., 2024; Chiang et al., 2023; Team et al., 2024a; Javaheripi et al., 2023; Almazrouei et al., 2023). These models were chosen due to their popularity, with over 100,000 downloads in the prior month on Hugging Face. [^3] As of November 2024 at https://huggingface.co/models The generated responses are obtained using top-k of 50 and a temperature of 0.75.
(a) Prompt template utilized to generate responses to BBH questions.
(b) Prompt template utilized to generate responses to SQuAD questions.
(c) Prompt template utilized to generate responses to the second turn and final of MT-Bench questions. To generate responses from the first turn, the template is identical to the BBH prompt.
Figure 1: Prompt templates utilized to generate responses from each of the varying benchmarks.
The model responses are then combined in a single prompt, as shown in Figure 2. To prevent judgment bias, the five responses to the respective benchmark question are included in the prompt in a random order as response options [A] - [E]. This random order was held constant across judgment replications, but shuffled for each new benchmark question judgment prompt. In addition to the five responses, the judgment prompt includes the original benchmark question (and context or previous turns if relevant) and prompts the LLM for a judgment of best response including Chain-of-Thought. Here, we utilize Chain-of-Thought from Wei et al. (2022) as it has been shown to dramatically improve LLM performance Feng et al. (2024); Suzgun et al. (2022); Chu et al. (2023). For the multi-turn responses from MT-Bench, options [A]-[E] were explicitly labeled as responses to the final turn, as shown in Figure 2(b).
(a) Prompt template utilized to generate judgments of responses to BBH and SQuAD questions.
(b) Prompt template utilized to generate judgments of responses to MT-Bench questions.
Figure 2: Prompt templates utilized to generate judgments of responses to each of the varying benchmarks.
Each LLM utilized for judgments was prompted 100 times, varying only the instantiation, to judge which of the responses to the given benchmark question is best out of the five responses provided. Given the employment of Chain-of-Thought in the initial responses to the question, a single best response should exist in response options [A] - [E] for a judge to uncover because each of the five responses contains reasoning in addition to the question response. Similar logic is employed in Zheng et al. (2023) in assessment of judges as well. The process is then repeated, varying the temperature of the judgments with temperatures of 1, 0.75, 0.5, 0.25, and 0 (machine epsilon) while holding top-k constant at 50. Temperature has long been known to impact the "novelty" of LLM outputs (Peeperkorn et al., 2024; Davis et al., 2024). Varying the temperature provides a window into the variability across instantiations and allows for assessment of the impact of fixed randomness.
@@
Selecting models that are high performers in the Chatbot Arena (Chiang et al., 2024), Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct are employed for the judgment task Zhu et al. (2023a); Team et al. (2024b); AI@Meta (2024). These selected models are all below the threshold of 10B parameters and are top performing models, ranked in the top 100, with rankings of 65 (Meta-Llama-3-8B-Instruct), 78 (Starling-LM-7B-beta), and 98 (Gemma-1.1-7b-it). In addition, we note that the selected models differ from those used for question answering in the prior step. This distinction is crucial to avoid circular reasoning, where LLMs would be judging responses they themselves generated. This approach also helps mitigate potential bias.
@@
Reliability of these judgments are then calculated using McDonald’s omega and employing the reliabiliPy package (Fernández, 2022). It is important to note that in calculating the reliability, judgments that did not provide a judgment of best response were treated as belonging to the same category, thereby resulting in the calculated reliability serving as an upper bound on reliability. This is further explored in the experiment results. Results are obtained using a pair of Quadro RTX 8000 GPUs with 48GB of memory each and the transformers library (Wolf et al., 2020).

### v2

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have experienced rapid advancements and widespread adoption, with applications ranging from marketing to biotechnology to music (Gozalo-Brizuela and Garrido-Merchán, 2023). As reliance upon LLMs grows, ensuring the quality and trustworthiness of their outputs becomes crucial. We often think about LLM quality in terms of accuracy (how often the LLM is correct). However, equally important are reliability, confidence, and calibration; these three distinct concepts are intertwined and essential for building genuine trust in LLM systems.
LLMs can express how certain, or ’confident’, they are about their responses. Calibration measures how well this confidence aligns with the LLM’s actual accuracy (Kotelanski et al., 2023; Spiess et al., 2024). A perfectly calibrated LLM that says it’s 90% confident should be correct about 9 out of 10 times. However, even a perfectly calibrated LLM can be unreliable.
Reliability refers to the consistency of an LLM’s outputs. A reliable weather forecasting LLM, for example, should give similar predictions for the same day, even if those predictions aren’t always accurate. An LLM that sometimes predicts sunshine and sometimes a downpour for the same day is unreliable, even if it’s well-calibrated and highly confident. This kind of consistency is what we aim to capture with internal consistency reliability. A reliable LLM can be consistently wrong, but it’s still reliable because its judgments are stable.
Why is reliability so important? Consider a scenario where an LLM is tasked with evaluating the quality of a medical imaging LLM used to diagnose tumors. Even if the LLM is perfectly calibrated (its confidence accurately reflects its diagnostic ability), if it gives different evaluations for the same AI on different runs, even with different, but accurate, confidence scores, doctors won’t know which evaluation to trust. The inconsistency makes the LLM unreliable and ultimately untrustworthy, regardless of its calibration and confidence.
@@
A singular output from an LLM is the result of a draw from a probabilistic distribution (Vaswani, 2017). Parameters that vary the resulting output, namely temperature and top-k, are increased to obtain more novel or "creative" results - a desirable quality of LLMs (Peeperkorn et al., 2024). However, this inherent stochastic nature raises concerns about individual outputs. Given that a single output from an LLM is a representation of only a single draw from the model’s distribution, trustworthiness of an LLM output is of utmost importance (Mittelstadt et al., 2023). Confidence and calibration are commonly used within NLP to understand this variability (Lin et al., 2023; Tian et al., 2023), but they don’t address reliability directly.
@@
Many research works aim to circumvent this issue by setting a fixed seed and using deterministic settings for the temperature and top-k parameters (Ouyang et al., 2023; Wei et al., 2024; Atil et al., 2024). These studies argue that if an LLM consistently produces the same output under these conditions, it can be considered reliable. However, consistent replication does not guarantee the reliability of the generated text.
Even with deterministic settings, a single LLM output remains a sample from the model’s probability distribution, subject to inherent randomness. This results in "fixed randomness," which can lead to significant limitations (Hellrich and Hahn, 2016). Consider our prior example of an LLM evaluator of a medical imaging LLM for tumor diagnosis. If the evaluating LLM, due to its own inherent randomness, consistently fails to identify a critical weakness in the target LLM in a single evaluation run, grave repercussions can arise. Eradicating fixed randomness is crucial to ensure model consistency and reduce the risk of these systematic errors. Internal consistency reliability addresses this problem by assessing how consistently the LLM applies its evaluation criteria across multiple evaluations, rather than relying on a single, potentially flawed output.
@@
While some research has focused on determining LLM reliability by comparing outputs to ground truth labels (Zhou et al., 2024; Fu et al., 2023a), this approach is often limited by the availability of accurate and comprehensive ground truth data. As LLMs are increasingly employed in complex and subjective domains, such as evaluating a medical imaging LLM for tumor diagnosis where even expert human diagnoses can vary, establishing a reliable ground truth becomes increasingly challenging. In current practice, LLM judgment models are often treated as raters, and inter-rater reliability is calculated on their outputs to assess reliability (Kollitsch et al., 2024; Wang et al., 2024). This metric, as we will show, is an insufficient quantification of reliability as the metric does not take into account the additional complexity of the LLM judge being a model itself and containing randomness.
The unique characteristics of LLMs as evaluators necessitate a more nuanced approach to assessing their reliability. We define reliability in this context as internal consistency reliability, a well-established statistical concept and the most common type of reliability employed across domains Henson (2001). While internal consistency is a well-established concept, its application to the LLM-as-a-judge paradigm, particularly with a focus on seed variation, has been largely overlooked. Internal consistency reliability measures how well different parts of an evaluation (individual LLM judgments) measure the same underlying construct (the LLM’s "true" judgment). Internal consistency reliability ensures that the LLM applies its evaluation criteria consistently across various inputs and conditions. Metrics like Cronbach’s alpha and McDonald’s omega provide quantifiable measures of this internal consistency, allowing us to assess the reliability of an LLM’s evaluations by examining the consistency of its responses across a range of prompts or tasks (Cronbach, 1951; McDonald, 2013; Malkewitz et al., 2023).
We introduce a novel framework that leverages McDonald’s omega to address a critical gap in current LLM evaluation methodologies by rigorously quantifying the stability and consistency of LLM judgments across multiple replicated evaluations of the same responses. Critically, this work pioneers the systematic investigation of the variability of an LLM evaluator’s judgments solely by varying the random seed, holding all other parameters constant. We assess how consistently LLMs apply their evaluation criteria across replications (i.e., multiple independent evaluations of the same responses with fixed parameters), investigating the impact of diverse question formats and varying difficulty levels using established benchmarks (BBH, SQuAD, MT-Bench). Our findings provide novel insights into the sensitivity of LLM judgments, revealing potential limitations in their reliability and highlighting the importance of considering this variability in high-stakes applications.  [^2] The code for the framework and analyses can be accessed at https://github.com/kaylaschroeder/llm_reliability.
@@ 2 Related Work
Internal consistency reliability itself, especially Cronbach’s alpha, has been heavily utilized as an assessor of question quality in LLM contexts (Shang et al., 2024; Chan et al., 2023b; Biri et al., 2023). Bhandari et al. (2024) uses Cronbach’s alpha to compare LLM-generated questions with human-authored ones; Shi et al. (2023) uses Cronbach’s alpha to assess the reliability of questions designed to evaluate LLMs. However, these studies focus on the reliability of questions, not the reliability of LLM judgments within the LLM-as-a-judge paradigm, a key distinction in our work.
@@
The concept of reliability appears widely in the LLM literature, but is not consistently defined. Some studies equate reliability with human-model agreement, requiring costly human evaluations (Sun et al., 2024; Zhang et al., 2024). Others rely on accuracy metrics, which necessitate ground truth labels (Zheng et al., 2024; Macherla et al., 2023; Wei et al., 2024). As the complexity and subjectivity of judgment tasks increase, obtaining reliable ground truth becomes increasingly challenging. Semantic consistency has also been proposed as a reliability measure (Govil et al., 2024; Raj et al., 2023), but this approach lacks a strong theoretical foundation and has potential pitfalls, as discussed in Schroeder and Wood-Doughty (2024). Many studies also use the term "reliability" without any rigorous definition (Tan et al., 2024; Liao et al., 2023; Gupta et al., 2024). This lack of a clear and consistent definition highlights the need for a more principled approach to measuring LLM judgment reliability.
@@
A common approach for establishing reliability across LLM judgment models is the employment of inter-rater reliability (Kollitsch et al., 2024; Wang et al., 2024). This metric, as we will show, is an insufficient quantification of reliability as the metric does not take into account the additional complexity of the LLM judge being a model itself and containing randomness. Beyond inter-rater reliability, the LLM-as-a-judge literature predominantly relies upon accuracy and bias metrics to assess reliability (Chen et al., 2024; Wei et al., 2024; Ye et al., 2024). Some researchers advocate for combining responses from different LLM judgment models to improve reliability (Gu et al., 2024; Patel et al., 2024). To quantify reliability of LLM judgments, human judgments are often used as a benchmark, with methods relying upon agreement with human evaluation and correlation with human-annotated error scores (Jung et al., 2024; Fu et al., 2023b; Dong et al., 2024). Unlike these approaches, which often rely on external benchmarks or human input, our work focuses on the internal consistency of LLM judgments, a critical aspect of reliability that has been largely overlooked in the LLM-as-a-judge paradigm. Our work pioneers the investigation of internal consistency reliability in this context, specifically by isolating the impact of random seed variation, addressing a significant gap in the literature.
3 Limitations of Current Reliability Measures
To illustrate the limitations of inter-rater reliability for LLM judgment reliability, we present a preliminary example. While inter-rater reliability, calculated as the percentage of agreement between raters and detailed further in (Hallgren, 2012), is a useful metric in some contexts, it is not well-suited for the unique characteristics of LLMs as judges (specifically, the inherent randomness within LLM evaluators themselves).
To demonstrate this, we performed an analysis using a simplified version of our Reliability Framework, which is described fully in Section 4. We obtained judgments from three LLM evaluators on responses to BIG-Bench Hard (BBH) questions from Suzgun et al. (2022), using a temperature of 0.25. Each LLM provided 100 judgments for each question, varying only the random seed. We then calculated the inter-rater reliability across the three LLM evaluators for each replication of judgments (one set per replication), for a total of 100 inter-rater reliability values.
Table 1 details the drastic variation in inter-rater reliability, ranging from 0.167 to 1.00. This disconcertingly wide range demonstrates that inter-rater reliability is highly sensitive to random seed variation, making it an unreliable indicator of true judgment quality. This instability in inter-rater reliability underscores the need for a more robust and appropriate measure of LLM judgment reliability, further motivating the development of our proposed (internal consistency) reliability framework.
Min Q1 Q2 Q3 Max
IRR 0.167 0.395 0.433 0.469 1.000
Table 1: Variation of the inter-rater reliability (IRR) across 3 LLM evaluators across replications.
4 Reliability Framework
Our framework evaluates the reliability of the LLM-as-a-judge paradigm across diverse question formats and difficulty levels. Using LLM responses to benchmark questions, judgment LLMs are prompted repeatedly to select the "best" response based on factors like accuracy, utility, and relevance. The LLM is prompted for evaluation 100 times, varying only the replication while holding other factors constant. All judgment results are then assessed to uncover reliability by applying McDonald’s omega.
The reliability framework begins by utilizing the widely accepted benchmarks BIG-Bench Hard (BBH) from Suzgun et al. (2022), SQuAD from Rajpurkar (2016), and MT-Bench from Zheng et al. (2023). These benchmarks offer two key advantages. First, they present diverse question formats: BBH and SQuAD questions are single-turn with specific target answers, while MT-Bench questions are multi-turn and open-ended. Second, they provide varying levels of difficulty, as evidenced by the significantly different accuracy rates of LLMs on BBH and SQuAD questions shown in Table 4. This diversity allows us to assess judgments under a range of conditions.
One question from each category within each benchmark is randomly selected, for a total of 55 questions. Next, five responses for each question are generated using Chain-of-Thought prompting, as illustrated in Figure 1. The variation in prompting by benchmark is a result of the structuring of each benchmark; BBH responses require a single question, SQuAD responses require a single question as well as context surrounding the question, and MT-Bench requires a two-turn dialogue.
(a) Prompt template for generating responses to BBH questions.
(b) Prompt template for generating responses to SQuAD questions.
(c) Prompt template for generating responses to the second turn and final of MT-Bench questions. To generate responses from the first turn, the template is identical to the BBH prompt.
Figure 1: Prompt templates to generate responses from each of the varying benchmark types.
The model responses are then combined in a single prompt, as shown in Figure 2. To prevent judgment bias, the five responses to the respective benchmark question are included in the prompt in a random order as response options [A] - [E]. This random order is held constant across judgment replications, but shuffled for each new benchmark question judgment prompt. In addition to the five responses, the judgment prompt includes the original benchmark question (and context or previous turns if relevant) and prompts the LLM for a judgment of best response including Chain-of-Thought. Here, Chain-of-Thought from Wei et al. (2022) is again utilized as it has been shown to dramatically improve LLM performance Feng et al. (2024); Suzgun et al. (2022); Chu et al. (2023). For the multi-turn responses from MT-Bench, options [A]-[E] are explicitly labeled as responses to the final turn, as shown in Figure 2(b).
(a) Prompt template for generating judgments of responses to BBH and SQuAD questions.
(b) Prompt template for generating judgments of responses to MT-Bench questions.
Figure 2: Prompt templates for generating LLM evaluations of responses for each of the varying benchmarks.
Each LLM utilized for judgments is prompted 100 times, varying only the replication, to evaluate which of the responses to a given benchmark question is best (as defined in the prompt) out of the five responses provided. Given the employment of Chain-of-Thought in the initial responses to a given benchmark question, a single best response should exist in response options [A] - [E] for a judge to uncover because each of the five responses contains reasoning in addition to the question response. Similar logic is employed in Zheng et al. (2023) in assessment of judgments as well.
@@ 4 Reliability Framework
The reliability of the judgments is then calculated using internal consistency reliability, a cornerstone of psychometric analysis and has been extensively studied. Cronbach’s alpha, introduced by Cronbach (1951), has been widely used to assess internal consistency, but relies on several assumptions that may not always hold (Malkewitz et al., 2023; Agbo, 2010). As proposed by McDonald (2013), McDonald’s omega addresses these limitations by accounting for a broader range of factor structures. Ultimately, omega provides a more robust measure of internal consistency and has been shown to be more robust to deviations from assumptions than alpha (Stensen and Lydersen, 2022). In this work, we adopt McDonald’s omega as our primary metric, recognizing its superior quantification of reliability.
@@ 4.1 McDonald’s Omega
where $n$ is the number of replications which, in this setting, is the number of times the LLM makes a judgment. The parameter $\lambda_{i}$ is the factor loading of the i-th replication. This represents the strength of the relationship between the i-th judgment and the underlying "true" judgment. In simpler terms, factor loadings indicate how much each replication contributes to our understanding of the overall judgment being measured. A higher factor loading suggests that the replication is a strong indicator of the underlying true judgment. The error variance of the i-th replication is given by $\theta_{ii}$, which represents the variability or "noise" associated with the i-th judgment. Finally, the covariance between the errors of replications i and j is given by $\theta_{ij}$. This measures the extent to which the errors in different judgments are related. Essentially, McDonald’s omega assesses how much of the observed variation in judgments reflects the true underlying value and how much is due to random error or inconsistencies in the measurement process.
@@
    Additive Measurement Error: Error in each LLM judgment replication is independent of the true judgment value. In simpler terms, the error does not systematically over- or underestimate the true judgment. This is a common psychometrics assumption, particularly when dealing with judgments that are expected to be relatively consistent (Alagumalai and Curtis, 2005; Wadkar et al., 2016; Liu et al., 2010).
@@
    Uncorrelated Errors: Errors in different LLM judgment replications are independent of each other. In other words, the error in one judgment does not influence the error in another judgment. This assumption is trivial in our setting, as each LLM judgment is generated independently so errors are guaranteed to be uncorrelated.
@@
    Single Latent Trait: All LLM judgments are attempting to measure the same underlying "true" judgment. In our case, each LLM judgment replication aims to identify the most accurate, relevant, and useful response to the same prompt, aligning with this assumption.
@@ 5 Experiments and Data
Responses to benchmark questions, using the templates outlined in Figure 1, are generated from five highly popular open-source LLMs: LLaMA-3-8B, Vicuna-7B-v1.5, Gemma-7B, Phi-2, and Falcon-7b (Dubey et al., 2024; Chiang et al., 2023; Team et al., 2024a; Javaheripi et al., 2023; Almazrouei et al., 2023). These models were chosen due to their popularity, with over 100,000 downloads in the prior month on Hugging Face. [^3] As of November 2024 at https://huggingface.co/models The generated responses are obtained using top-k of 50 and a temperature of 0.75.
@@
Selecting models that are high performers in the Chatbot Arena (Chiang et al., 2024), Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct are employed for the judgment task Zhu et al. (2023); Team et al. (2024b); AI@Meta (2024). These selected models are all below the threshold of 10B parameters and are top performing models, ranked in the top 100, with rankings of 65 (Meta-Llama-3-8B-Instruct), 78 (Starling-LM-7B-beta), and 98 (Gemma-1.1-7b-it). In addition, we note that the selected models differ from those used for question answering in the prior step. This distinction is crucial to avoid circular reasoning, where LLMs would be judging responses they themselves generated. This approach also helps mitigate potential bias.
To further investigate the influence of fixed randomness on LLM outputs, we systematically varied the temperature parameter during the LLM evaluation process. Temperature, a key hyperparameter in LLMs, significantly impacts the randomness (or "novelty") of their outputs (Peeperkorn et al., 2024; Davis et al., 2024). We evaluated LLM judgments at five temperature levels: 1.0, 0.75, 0.5, 0.25, and 0 (machine epsilon), while keeping the top-k sampling parameter constant at 50. This systematic variation in temperature allowed us to observe how changes in the level of randomness within the LLM’s generation process affect the variability and reliability of the subsequent judgments, providing a window into the impact of fixed randomness.
@@
The calculation of McDonald’s omega employs the reliabiliPy package (Fernández, 2022). It is important to note that in calculating the reliability, all judgments that did not provide a judgment of best response were treated as belonging to the same category ("Non Response") which was included as a category in addition to the five response categories "A" through "E". Consequently, the resulting calculated reliability serves as an upper bound on reliability because the variability in non-response judgments is ignored. This is further explored in the experiment results. Results are obtained using a pair of Quadro RTX 8000 GPUs with 48GB of memory each and the transformers library (Wolf et al., 2020).